makenzieee:

Susan is a college student with two part-time jobs. She earns $10 per hour tutoring elementary students in math. She earns $15 per hour cleaning in the library. Her goal is to earn at least $240 per week, but because of college, she does not work more than 20 hours each week. PART B Which combinations allow Susan to work no more than 20 hours in one week and earn at least $240? Select the three correct combinations A 9 hours tutoring and 10 hours cleaning the library B 10 hours tutoring and 10 hours cleaning the library C 11 hours tutoring and 8 hours cleaning the library D 12 hours tutoring and 8 hours cleaning the library E 13 hours tutoring and 7 hours cleaning the library F 14 hours tutoring and 7 hours cleaning the library

jhonyy9:

hi ! any idea ?

jhonyy9:

try these choices step by step and see on what will get in the time of 20 hours this sum of $240

jhonyy9:

for example A. 9 h tutoring and 10 h cleaning 9*10 = 90 10*15 = 150 so 90 + 150 = 240 what is your opinion about this choice ? so in this way try these others

jhonyy9:

do you understand it ?

jhonyy9:

hi ! do you have again questions ?